A watermelon is suspended from a spring and causes the spring to expand 0.54 m. If the spring constant is 131 N/m, what is the mass of the water melon?
1
Expert's answer
2020-12-02T07:15:26-0500

We can find the mass of the watermelon from the Hooke's Law:


"F=kx,""mg=kx,""m=\dfrac{kx}{g}=\dfrac{131\ \dfrac{N}{m}\cdot 0.54\ m}{9.8\ \dfrac{m}{s^2}}=7.2\ kg."

Answer:

"m=7.2\ kg."